# Diffusion posterior sampling for simulation-based inference in tall data settings

### 11 Apr 2024 v1

@@ Appendix B Positive Definiteness of $\Lambda(\theta)$
  • For $t\to 1$: the forward kernel and diffused distribution approach the noise, i.e. $p_{t}(\theta\mid x)\underset{t\to 1}{\to}\mathcal{N}(\theta;0,\mathbf{I}_{m})$ and $q_{t\mid 0}(\theta\mid\theta_{0})\underset{t\to 1}{\to}\mathcal{N}(\theta;0,\mathbf{I}_{m})$. The backward kernel is thus equivalent to the target distribution:
@@ D.1 Gaussian case
Replacing the score from (26) in the formulas from (Boys et al.,, 2023), we get the following expressions for the mean and covariance matrix of the posterior backward diffusion kernel $p_{0\mid t}(\theta_{0}\mid\theta,x)$:
@@
We now consider the tall data setting with $n$ i.i.d. observations $x^{\star}_{1:n}=\{x^{\star}_{1},\dots,x^{\star}_{n}\}$ and compute the correction term $\nabla_{\theta}\ell_{\lambda}(\theta,x^{\star}_{1:n})$. Note that the covariance matrices above are independent of $\theta$ and therefore $\Lambda(\theta)$ does not depend on $\theta$ and is referred to as $\Lambda$. Consequently, we only need to consider the terms that depend on the means $\mathrm{\mu}_{t}(\theta,x^{\star}_{j})$ or $\mathrm{\mu}_{t,\lambda}(\theta)$ when computing the gradient w.r.t. $\theta$.
From Boys et al., (2023) we have
$\nabla\zeta_{0|t}(\theta,x)=-\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\mu_{t}(\theta,x)\,,\quad\nabla\zeta_{0|t,\lambda}(\theta)=-\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\mu_{t,\lambda}(\theta)\,.$ (27)
For $\zeta_{0|t,\mathrm{all}}$, write
$\eta_{0|t}(\theta)$ $=\sum_{j=1}^{N}\Sigma_{t}^{-1}\left[\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\Sigma_{t}\theta+\upsilon_{t}\Big{(}\alpha_{t}\Sigma_{\mathrm{post}}+\upsilon_{t}\mathbf{I}_{m}\Big{)}^{-1}\mu_{\mathrm{post}}(x^{\star}_{j})\right]$
$+(1-n)\Sigma_{t,\lambda}^{-1}\left[\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\Sigma_{t,\lambda}\theta+\upsilon_{t}\Big{(}\alpha_{t}\Sigma_{\lambda}+\upsilon_{t}\mathbf{I}_{m}\Big{)}^{-1}\mu_{\lambda}\right]$
$=\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\theta+\upsilon_{t}\Sigma_{t}^{-1}\Big{(}\alpha_{t}\Sigma_{\mathrm{post}}+\upsilon_{t}\mathbf{I}_{m}\Big{)}^{-1}\sum_{j=1}^{N}\mu_{\mathrm{post}}(x^{\star}_{j})+(1-n)\upsilon_{t}\Sigma_{t,\lambda}^{-1}\Big{(}\alpha_{t}\Sigma_{\lambda}+\upsilon_{t}\mathbf{I}_{m}\Big{)}^{-1}\mu_{\lambda}\,.$
We need to assure the positive definiteness of $\Lambda(\theta)$. Let $v\in\mathbb{R}^{d}$ such that $\|v\|=1$. Thus,
$v^{\top}\Lambda(\theta)v\propto\frac{1}{\upsilon_{t}}+\alpha^{2}_{t}(1-n)v^{\top}\Sigma_{\lambda}v+n\alpha^{2}_{t}v^{\top}\Sigma_{\rm{post}}v\,.$ (28)
Therefore, for $\Lambda(\theta)$ to be positive definite is equivalent to
$v^{\top}\Sigma_{\rm{post}}v>\frac{n-1}{n}v^{\top}\Sigma_{\lambda}v-\frac{1}{n\upsilon_{t}\alpha^{2}_{t}}\,,$ (29)
for all $v\in\mathbb{R}^{d}$ such that $\|v\|=1$. Note that it is not trivial to find meaningful worst case scenario sufficient condition for the inequality above. To see why, let $e_{\rm{post}}^{\min}$ be the smallest eigenvalue of $\Sigma_{\rm{post}}$ and $e_{\lambda}^{\max}$ the biggest eigenvalue of $\Sigma_{\lambda}$. Then, a sufficient condition would be $e_{\lambda}^{\max}<e_{\rm{post}}^{\min}$, since
$v^{\top}\Sigma_{\rm{post}}v>e_{\rm{post}}^{\min}>e_{\lambda}^{\max}>\frac{n-1}{n}v^{\top}\Sigma_{\lambda}v-\frac{1}{n\upsilon_{t}\alpha^{2}_{t}}\,.$ (30)
But this does not make much sense, since we are asking the prior to be more concentrated than the posterior! Therefore, we obtain $\nabla\zeta_{0|t,\mathrm{all}}=-\sqrt{\alpha_{t}}\Lambda^{-1}\eta_{0|t}(\theta)/\upsilon_{t}$, which yields
$\nabla\ell_{\lambda}(\theta,x^{\star}_{1:n})=-\frac{\sqrt{\alpha_{t}}}{\upsilon_{t}}\left(\sum_{j=1}^{n}\mu_{t}(\theta,x^{\star}_{j})+(1-n)\mu_{t,\lambda}(\theta)-\Lambda^{-1}\eta_{0|t}(\theta)\right)\,.$ (31)
D.2 Score of the diffused Uniform prior
@@ Appendix E Implementation details
To evaluate the accuracy of the sampling algorithms, we compute the sliced Wasserstein (sW) distance using the implementation form POT with 1 00010001\,0001 000 projections and the Maximum Mean Discrepancy (MMD) from the sbi package. We also evaluate the distance to the true parameters $\theta^{\star}$ used to simulated the observations $x^{\star}_{1:n}$ by computing the MMD to the Dirac function $\delta_{\theta^{\star}}$ (MMD to Dirac).
@@ Appendix F Loss functions, training strategy and model selection
Figure 7: Loss functions for the Jansen & Rit Neural Mass Model. The validation loss was computed on a held-out validation set of $20\%$ of the training dataset of size $N_{\mathrm{train}}=50~{}000$.
@@ Appendix G Additional results for the toy models
Table 3: Sliced Wasserstein (sW) and total elapsed time for the Gaussian Mixture toy problem with $d=10$, $n=32$ and $\epsilon=10^{-2}$ per algorithm and number of generation steps. Mean and std over 5 different seeds.
Figure 8: Normalized sliced Wasserstein (sW) distance as a function of the number of observations $n$ for GAUSS, JAC and LANGEVIN with different levels of $\epsilon$. Results are shown for the Gaussian example in several dimensions $m\in[2,4,8,10,16,32]$. Mean and std over 5 different seeds.
Figure 9: Sliced Wassersteiin (sW) distance as a function of the number of observations $n$ for the different methods with different levels of $\epsilon$. Results are shown for the Gaussian example in several dimensions $m\in[2,4,8,10,16,32]$. Mean and std over 5 different seeds.
@@ Appendix H Additional results for SBI benchmarks
H.1 Sliced Wasserstein (sW) distance)
@@
H.2 Maximum Mean Discrepancy (MMD)
As a complementary result to the one reported with the sliced Wassersetin distance, the MMD metric highlights an interesting point: Figures 12 and 13 show a general trend of increasing MMD values for higher $n$. This can be explained by the accumulation of approximation errors, as we sum over $n$ evaluations of the score model to obtained the tall posterior score conditioned on $n$ observations. We refer the reader to Appendix K, which investigates a possible solution to this issue: partially factorized score-based posterior sampling algorithms, such as PF-NPSE proposed by (Geffner et al.,, 2023).
@@ H.2 Maximum Mean Discrepancy (MMD)
H.3 Distance to the Dirac of the true parameters (MMD to Dirac)
@@
Figure 14: MMD as a function of $N_{\mathrm{train}}\in[10^{3},3.10^{3},10^{4},3.10^{4}]$ between the marginals of the approximate tall posterior distribution $p(\theta\mid x^{\star}_{1,n})$ and the Dirac of the true parameters $\theta^{\star}$ used to simulate the observations $x^{\star}_{1,n}$ (for $n\in[1,8,14,22,30]$). Mean and std over 25 different parameters $\theta^{\star}\sim\lambda(\theta)$.
@@
Figure 15: MMD as a function of $n\in[1,8,14,22,30]$ between the marginals of the approximate tall posterior $p(\theta\mid x^{\star}_{1,n})$ (for $N_{\mathrm{train}}\in[10^{3},3.10^{3},10^{4},3.10^{4}]$) and the Dirac of the true parameters $\theta^{\star}$ used to simulate the observations $x^{\star}_{1,n}$. Mean and std over 25 different parameters $\theta^{\star}\sim\lambda(\theta)$.
Appendix I Results for additional SBI Benchmark examples
@@
To complete our empirical study, we added results for additional examples from the SBI benchmark (Lueckmann et al., 2021b, ): Gaussian Linear, GMM, GMM (uniform)101010same as GMM but with a Uniform prior., B-GLM/ (raw) [^11] Bernoulli GLM with summary statics / high dimensional raw data. and Two Moons. These new results allows us to compare the performance of our proposal on other challenging situations, such as when scaling to highly structured (e.g. multimodal) posteriors and high-dimensional observation spaces. Note that these examples go a step further as compared to the experiments carried out by Geffner et al., (2023), including non-Gaussian priors [^12] Note that this was already the case for SLCP with Uniform prior.. Figures 16 and 17 respectively report the sW and MMD as a function of $N_{\mathrm{train}}$.
Figure 16: Results for additional benchmark examples (sW as a function of $N_{\mathrm{train}}$).
Figure 17: Results for additional benchmark examples (MMD as a function of $N_{\mathrm{train}}$).
Appendix J Limitations of Langevin sampling
@@
Appendix K Extensions: Classifier-free guidance and Partial Factorization
@@
Classifier-free guidance (CFG).
It is possible to implicitly learn the prior score via the classifier-free guidance approach (Ho and Salimans,, 2022), which essentially consists in randomly dropping the context variables when training the posterior score model (e.g. $20\%$ of the time). This is useful in cases where the diffused prior score cannot be computed analytically. Figure 19 displays the sliced Wasserstein and MMD as a function of $N_{\mathrm{train}}$ and compares the results obtained when using our proposition with the learned vs. the analytical prior score (resp. GAUSS (CFG) vs. GAUSS). The results are highly accurate for the Log-Normal priors of Lotka-Volterra and SIR, but less satisfying for the Uniform prior in SLCP. We think that this is caused by the discontinuities of the Uniform distribution. In summary, it seems that (under some smoothness assumptions) it is possible to learn the prior score via the classifier-free guidance approach.
(a) Sliced Wasserstein (sW).
(a) Sliced Wasserstein (sW).
(b) Maximum Mean Discreapancy (MMD).
Figure 19: Comparison between the results obtained with our algorithm when combined with the analytical prior score and the one learned via classifier-free guidance (resp. GAUSS and GAUSS (CFG)). The plots show the sW and MMD w.r.t. the true tall posterior $p(\theta\mid x^{\star}_{1:n})$ as a function of $N_{\mathrm{train}}\in\{10^{3},3.10^{3},10^{4},3.10^{4}\}$ and for $n\in\{1,8,14,22,30\}$.
Partial factorization (PF-NPSE).
@@
We investigated the performance of PF-NPSE on the SBI benchmarks (Lotka-Volterra, SIR and SLCP). For each of the three examples we trained a PF-NPSE model targeting the score models for the law of $\theta$ given $x_{1:n_{\mathrm{max}}}$ for $n_{\mathrm{max}}\in\{1,3,6,30\}$. Figure 20 displays the Sliced Wasserstein (sW) for $n=30$ observations as a function of the $N_{\mathrm{train}}$ for samples obtained with the partially factorized LANGEVIN and GAUSS samplers and for all $n_{\mathrm{max}}$. The extreme case $n_{\mathrm{max}}=1$ corresponds to the original ”fully” factorized version of the samplers. $n_{\mathrm{max}}=n=30$ correspond to the other extreme case with no factorization, but maximum simulation budget. We can see that the optimal sW values lie in the middle of the spectrum (i.e. for $n_{\mathrm{max}}=3,6$), which corresponds to what was concluded in (Geffner et al.,, 2023). Note that the performance of LANGEVIN is drastically improved for $n_{\mathrm{max}}>1$, while GAUSS all results are close. In any case, the results suggests that a practitioner will gain in choosing (a small enough) $n_{\mathrm{max}}>1$.
Figure 20: Results obtained with the partially factorized LANGEVIN and GAUSS samplers to infer the tall posterior conditioned on a total number of observations $n=30$, for $n_{\mathrm{max}}=1,3,6,30$. We report the sliced Wasserstein (sW) and MMD w.r.t. the true tall posterior $p(\theta\mid x^{\star}_{1:n})$ as a function of $N_{\mathrm{train}}\in\{10^{3},3.10^{3},10^{4},3.10^{4}\}$. The extreme case $n_{\mathrm{max}}=1$ corresponds to the original ”fully” factorized posterior samplers. $n_{\mathrm{max}}=n=30$ correspond to the case with no factorization, but maximum simulation budget. We can see that the optimal distance values lie in the middle of this spectrum (i.e. for $n_{\mathrm{max}}=3,6$).

### 07 Jun 2024 v2

@@ Appendix B Positive Definiteness of $\Lambda(\theta)$
  • For $t\to 1$: the forward kernel and the diffused data distribution approach the noise distribution: $p_{t}(\theta\mid x)\underset{t\to 1}{\to}\mathcal{N}(\theta;0,\mathbf{I}_{m})$ and $q_{t\mid 0}(\theta\mid\theta_{0})\underset{t\to 1}{\to}\mathcal{N}(\theta;0,\mathbf{I}_{m})$. The backward kernel is thus equivalent to the target data distribution:
@@ D.1 Gaussian case
Replacing the score from (26) in the formulas from (Boys et al.,, 2023), we get the following expressions for the mean and covariance matrix of the backward diffusion kernel $p_{0\mid t}(\theta_{0}\mid\theta,x)$:
@@
D.2 Score of the diffused Log-Normal prior
The Log-Normal distribution can easily be transformed into a Gaussian distribution:
$\theta\sim\mathrm{LogNormal}(m,s)\Rightarrow\log\theta\sim\mathcal{N}(m,s)\,.$
We can therefore directly apply the Gaussian case from the previous section, after transforming the data into the log-space.
D.3 Score of the diffused Uniform prior
@@ Appendix E Implementation details
To evaluate the accuracy of the sampling algorithms on the SBI benchmark examples, we compute the distance between 10 0001000010\,00010 000 samples of the true and estimated posterior using three different metrics: the sliced Wasserstein (sW) distance from the POT python package with 1 00010001\,0001 000 projections, the Maximum Mean Discrepancy (MMD) and the Classifier-Two-Sample Test (C2ST), both from the sbibm package, with default parameters. For the Jansen and Rit Neural Mass Model, our real SBI example from computational neuroscience, no samples from the true posterior are available and the above metrics cannot be computed. We therefore use the local Classifier-Two-Sample Test ($\ell$-C2ST) recently proposed by Linhart et al., (2023), as implemented in the sbi toolbox , with default parameters.
Code and compute resources.
The code and guidelines to reproduce all numerical experiments is provided in the following github repository: https://github.com/JuliaLinhart/diffusions-for-sbi. This includes specific information about the code environment and installation requirements, mentioned in the ”readme.md” file. In this url, we also provide pre-generated data and pre-computed results that are necessary to quickly generate all figures of the paper without extensive computation time.
@@ Appendix F Loss functions, training strategy and model selection
Figure 6: Loss functions for the Jansen & Rit Neural Mass Model obtained for score networks trained over 5 000 epochs with the Adam optimzer, a batch size of $256$ and learning rates $\mathrm{lr}=1e^{-3}$ (orange) and $\mathrm{lr}=1e^{-4}$ (blue). The validation loss was computed on a held-out validation set of $20\%$ of the training dataset of size $N_{\mathrm{train}}=50~{}000$. Dashed lines indicate the epoch at early stopping time. The chosen model corresponds to the smallest validation loss, here obtained for a learning rate of $1e^{-4}$.
@@ Appendix G Additional results for the toy models
Table 3: Sliced Wasserstein (sW) distance (normalized) and total elapsed time for the Gaussian Mixture toy problem with $d=10$, $n=32$ and $\epsilon=10^{-2}$ per algorithm and number of generation steps. Mean and std over 5 different seeds.
Figure 7: Sliced Wasserstein (sW) distance (normalized) as a function of the number of observations $n$ for GAUSS, JAC and LANGEVIN with different levels of $\epsilon$. Results are shown for the Gaussian example in several dimensions $m\in[2,4,8,10,16,32]$. Mean and std over 5 different seeds.
Figure 8: Sliced Wasserstein (sW) distance (normalized) as a function of the number of observations $n$ for the different methods with different levels of $\epsilon$. Results are shown for the Gaussian example in several dimensions $m\in[2,4,8,10,16,32]$. Mean and std over 5 different seeds.
@@ Appendix H Additional results for SBI benchmarks
We here include more detailed results for the three SBI benchmark tasks from Section 4.2, i.e. for more values of $n$ and additional evaluation metrics. As a complement to the sliced Wassersetin (sW) distance, we report results obtained for the Maximum Mean Discrepancy (MMD) and the Classifier-Two-Sample Test (C2ST) metrics from the sbibm package, respectively in Appendix H.2 and H.3. These metrics are less discriminative but highlight an interesting point: Figures 12 and 14 show increasing MMD and C2ST values for higher $n$. This can be explained by the accumulation of approximation errors, as we sum over $n$ evaluations of the score model to obtained the tall posterior score for $n$ observations. We refer the reader to Appendix L.2, which investigates a possible solution to this issue: partially factorized score-based posterior sampling algorithms, such as PF-NPSE proposed by (Geffner et al.,, 2023).
H.1 Sliced Wasserstein distance
@@
H.2 Maximum Mean Discrepancy
@@
H.3 Classifier-Two-Sample Test
@@
Figure 13: C2ST as a function of $N_{\mathrm{train}}\in[10^{3},3.10^{3},10^{4},3.10^{4}]$ between between the samples obtained by each algorithm and the true tall posterior distribution $p(\theta\mid x^{\star}_{1,n})$ (for $n\in[1,8,14,22,30]$). Mean and std over 25 different parameters $\theta^{\star}\sim\lambda(\theta)$.
@@
Figure 14: C2ST as a function of $n\in[1,8,14,22,30]$ between the samples obtained by each algorithm and the true tall posterior distribution $p(\theta\mid x^{\star}_{1,n})$ (for $N_{\mathrm{train}}\in[10^{3},3.10^{3},10^{4},3.10^{4}]$) and the Dirac of the true parameters $\theta^{\star}$ used to simulate the observations $x^{\star}_{1,n}$. Mean and std over 25 different parameters $\theta^{\star}\sim\lambda(\theta)$.
Appendix I Results for additional tasks from the SBI benchmark
@@
To complete our empirical study, we added results for additional examples from the SBI benchmark (Lueckmann et al.,, 2021): Gaussian Linear, GMM, GMM (uniform)777same as GMM but with a Uniform prior., B-GLM/ (raw) [^8] Bernoulli GLM with summary statics / high dimensional raw data. and Two Moons. These new results allows us to compare the performance of our proposal on other challenging situations, such as when scaling to highly structured (e.g. multimodal) posteriors and high-dimensional observation spaces. Note that these examples go a step further as compared to the experiments carried out by Geffner et al., (2023), including non-Gaussian priors [^9] Note that this was already the case for SLCP with Uniform prior.. Figures 15, 16 and 17 respectively report the sW, MMD and C2ST as a function of $N_{\mathrm{train}}$. They all show that our algorithm outperforms the Langevin sampler, with smaller distance values everywhere but for the GMM (uniform) example. Interestingly, the certainly very discriminative C2ST metric, shows that all three algorithms fail to infer the tall posterior for the Two Moons and B-GLM/ (raw) examples.
Figure 15: Sliced Wasserstein (sW) distance as a function of $N_{\mathrm{train}}\in[10^{3},3.10^{3},10^{4},3.10^{4}]$ between between the samples obtained by each algorithm and the true tall posterior distribution $p(\theta\mid x^{\star}_{1,n})$ (for $n\in[1,8,14,22,30]$). Mean and std over 25 different parameters $\theta^{\star}\sim\lambda(\theta)$.
Figure 16: MMD as a function of $N_{\mathrm{train}}\in[10^{3},3.10^{3},10^{4},3.10^{4}]$ between between the samples obtained by each algorithm and the true tall posterior distribution $p(\theta\mid x^{\star}_{1,n})$ (for $n\in[1,8,14,22,30]$). Mean and std over 25 different parameters $\theta^{\star}\sim\lambda(\theta)$.
Figure 17: C2ST as a function of $N_{\mathrm{train}}\in[10^{3},3.10^{3},10^{4},3.10^{4}]$ between between the samples obtained by each algorithm and the true tall posterior distribution $p(\theta\mid x^{\star}_{1,n})$ (for $n\in[1,8,14,22,30]$). Mean and std over 25 different parameters $\theta^{\star}\sim\lambda(\theta)$.
Appendix J Results for the Jansen-Rit Neural Mass Model
In this section, we report results obtained for the simplified version (3D) and the full (4D) Jansen and Rit Neural Mass Model (JRNMM).
In both cases, we first evaluate the accuracy of each posterior sampling algorithms (GAUSS, JAC and LANGEVIN) w.r.t. the true tall posterior, using the local Classifier-Two-sample Test ($\ell$-C2ST) diagnostic proposed by Linhart et al., (2023) and implemented in the sbi toolbox. $\ell$-C2ST is a frequentist hypothesis test that evaluates the null hypothesis of sample equality between two conditional distributions by training a classifier on data from the respective joint distributions. The test statistic takes values in $[0,0.25]$ and is defined by the mean squared error (MSE) between the predicted class probabilities and one half, i.e. the chance level where the classifier is unable to distinguish between the two data classes. Higher MSE values indicate more distributional differences.
Figures 18 and 19 respectively show the $\ell$-C2ST results for the 3D and 4D JRNMM cases, obtained using the MLPClassifier from scikit-learn, with default parameters from the sbi implementation and trained using 10 0001000010\,00010 000 samples $(\Theta_{i},X_{i,1:n})$ from the joint distributions corresponding to the estimated and true tall data posteriors. We report the mean and std over 5 different seeds used for the initialization of the classifier. On the left, we plot the $\ell$-C2ST statistics, computed on 10 0001000010\,00010 000 samples of the approximate posterior obtained by each sampling algorithms for a given observation set $x^{\star}_{1:n}\sim p($. On the right, we display the associated p-values computed by comparing the obtained statistics to the null hypothesis, estimated over $100$ trials. In both cases, the GAUSS algorithm consistently yields test statistics close to 00 and p-values above the significance level set at $\alpha=0.05$, which means that there is no evidence that the estimated posterior is not statistically consistent with the true tall posterior at $x^{\star}_{1:n}$. This is not the case for JAC, whose un-clipped version yields MSE values outside the plot limits and LANGEVIN, for which the test is more easily rejected in the 3D case (large std on pvalues that overlap with the significance level) and very clearly rejected in the 4D case for $n>14$ (p-value below the significance level).
In a second time, we evaluate the the ability of our tall posterior to concentrate around the true parameters $\theta^{\star}$ generating observations $x_{1},\dots,x_{n}\sim p(x\mid\theta^{\star})$ as $n$ increases. This is quantified using the Maximum Mean Discrepancy (MMD) between the posterior marginals and the Dirac distribution $\delta_{\theta^{\star}}$, computed on 10 0001000010\,00010 000 samples obtained with GAUSS.
Figures 18 and 19 respectively show the results for the 3D and 4D JRNMM cases. On the left, we plot the MMD as a function of the number $n$ of observations. On the right are displayed the marginals corresponding to the posterior samples obtained with GAUSS for observation sets $x^{\star}_{1:n}$ of increasing size. For the 3D case we observe, a progressive concentration around $\theta^{\star}$, with sharper posterior marginals and decreasing MMD. For the 4D JRNMM, the tall posterior takes a “sharpened banana” shape: while it is able to concentrate around $(C,g)$, the dispersion along the dimensions of $(\mu,\sigma)$ is sustained.
(a) Accuracy of the sampling algorithms w.r.t. the true tall posterior. We show $\ell$-C2ST statistics (left) and corresponding p-values (right) computed for samples obtained with GAUSS, JAC and LANGEVIN at $x^{\star}_{1:n}$ for $n\in[1,8,14,22,30]$. Mean and std over 5 seeds.
(b) Concentration of the tall posterior. Left: MMD between the marginals of the approximate posterior obtained with GAUSS and the Dirac of the true parameters $\theta^{\star}=(125,220,2000)$ (black dashed lines) used to simulate the observations $x^{\star}_{1:n}$. Right: Histograms of the 1D marginals of the posterior samples obtained with GAUSS for $30$ single observations $x^{\star}_{1},\dots,x^{\star}_{30}$ ($n=1$) and for observation sets $x^{\star}_{1:n}$ of increasing size.
Figure 18: Inference on the 3D JRNMM (fixed $g=0$).
(a) Accuracy of the sampling algorithms w.r.t. the true tall posterior. We show $\ell$-C2ST statistics (left) and corresponding p-values (right) computed for GAUSS, JAC and LANGEVIN at $x^{\star}_{1:n}$ for $n\in[1,8,14,22,30]$. Mean and std over 5 seeds.
(b) Concentration of the tall posterior. Left: MMD between the marginals of the approximate posterior obtained with GAUSS and the Dirac of the true parameters $\theta^{\star}=(125,220,2000,0)$ (black dashed lines) used to simulate the observations $x^{\star}_{1:n}$. Right: Histograms of the 1D and 2D marginals of the posterior samples obtained with GAUSS for a single observations $x^{\star}$ ($n=1$) and for observation sets $x^{\star}_{1:n}$ of increasing size.
Figure 19: Inference on the full (4D) JRNMM.
Appendix K Limitations of Langevin sampling
@@
Appendix L Beyond the proposal: Classifier-free guidance and Partial Factorization
@@
L.1 Classifier-free guidance (CFG).
It is possible to implicitly learn the prior score via the classifier-free guidance approach (Ho and Salimans,, 2021), which essentially consists in randomly dropping the context variables when training the posterior score model (e.g. $20\%$ of the time). This is useful in cases where the diffused prior score cannot be computed analytically. Figure 21 displays the sliced Wasserstein (sW) distance as a function of $N_{\mathrm{train}}$ and compares the results obtained when using our proposition with the learned vs. the analytical prior score (resp. GAUSS (CFG) vs. GAUSS). We also report the results obtained Maximum Mean Discrepancy (MMD) and Classifier-Two-Sample Test (C2ST) accuracy respectevely in Figures 22 and 23. The results are highly accurate for the Log-Normal priors of Lotka-Volterra and SIR, but less satisfying for the Uniform prior in SLCP. We think that this is caused by the discontinuities of the Uniform distribution. In summary, it seems that (under some smoothness assumptions) it is indeed possible to learn the prior score via the classifier-free guidance approach.
Figure 21: Sliced Wasserstein (sW) distance w.r.t. the true tall posterior $p(\theta\mid x^{\star}_{1:n})$ as a function of $N_{\mathrm{train}}\in\{10^{3},3.10^{3},10^{4},3.10^{4}\}$ and for $n\in\{1,8,14,22,30\}$. Comparison between the results obtained with our algorithm when combined with the analytical prior score and the one learned via classifier-free guidance (resp. GAUSS and GAUSS (CFG)).
Figure 22: Maximum Mean Discreapancy (MMD) w.r.t. the true tall posterior $p(\theta\mid x^{\star}_{1:n})$ as a function of $N_{\mathrm{train}}\in\{10^{3},3.10^{3},10^{4},3.10^{4}\}$ and for $n\in\{1,8,14,22,30\}$. Comparison between the results obtained with our algorithm when combined with the analytical prior score and the one learned via classifier-free guidance (resp. GAUSS and GAUSS (CFG)).
Figure 23: Classifier-Two-Sample Test (C2ST) accuracy w.r.t. the true tall posterior $p(\theta\mid x^{\star}_{1:n})$ as a function of $N_{\mathrm{train}}\in\{10^{3},3.10^{3},10^{4},3.10^{4}\}$ and for $n\in\{1,8,14,22,30\}$. Comparison between the results obtained with our algorithm when combined with the analytical prior score and the one learned via classifier-free guidance (resp. GAUSS and GAUSS (CFG)).
L.2 Partial factorization (PF-NPSE).
@@
We investigated the performance of PF-NPSE on the SBI benchmarks (Lotka-Volterra, SIR and SLCP). For each of the three examples we trained a PF-NPSE model targeting the score models for the law of $\theta$ given $x_{1:n_{\mathrm{max}}}$ for $n_{\mathrm{max}}\in\{1,3,6,30\}$. Figure 24 displays the sliced Wasserstein (sW), Maximum Mean Discrepancy (MMD) and the Classifier-Two-Sample Test (C2ST) accuracy for $n=30$ observations as a function of the $N_{\mathrm{train}}$ for samples obtained with the partially factorized LANGEVIN and GAUSS samplers and for all $n_{\mathrm{max}}$. The extreme case $n_{\mathrm{max}}=1$ corresponds to the original ”fully” factorized version of the samplers. $n_{\mathrm{max}}=n=30$ correspond to the other extreme case with no factorization, but maximum simulation budget. We can see that the optimal sW values lie in the middle of the spectrum (i.e. for $n_{\mathrm{max}}=3,6$), which corresponds to what was concluded in (Geffner et al.,, 2023). Note that the performance of LANGEVIN is drastically improved for $n_{\mathrm{max}}>1$, while GAUSS all results are close. In any case, the results suggests that a practitioner will gain in choosing (a small enough) $n_{\mathrm{max}}>1$.
Figure 24: Results obtained with the partially factorized LANGEVIN and GAUSS samplers to infer the tall posterior conditioned on a total number of observations $n=30$, for $n_{\mathrm{max}}=1,3,6,30$. We report the sliced Wasserstein (sW), Maximum Mean Discrepancy (MMD) and the Classifier-Two-Sample Test (C2ST) accuracy w.r.t. the true tall posterior $p(\theta\mid x^{\star}_{1:n})$ as a function of $N_{\mathrm{train}}\in\{10^{3},3.10^{3},10^{4},3.10^{4}\}$. The extreme case $n_{\mathrm{max}}=1$ corresponds to the original ”fully” factorized version of the samplers. $n_{\mathrm{max}}=n=30$ correspond to the other extreme case with no factorization, but maximum simulation budget. We can see that the optimal distance values lie in the middle of this spectrum (i.e. for $n_{\mathrm{max}}=3,6$).